OpenStudy (jaypeeheat):

Two consecutive whole numbers that square root of 91 lies between?

OpenStudy (mathstudent55):

Complete the following lines: 1^2 = 1*1 = 1 2^2 = 2*2 = 4 3^2 = 3*3 = 9 4^2 = 4*4 = 16 5^2 = 5*5 = 25 6^2 = 7^2 = 8^2 = 9^2 = 10^2 = 11^2 =